OpenStudy (anonymous):

Solve for x: -4(x-3)=16

OpenStudy (anonymous):

-4(x-3)=16 distribute the -4 on the left hand side first

OpenStudy (anonymous):

What do you get after doing that?

OpenStudy (anonymous):

I got -7, but someone else got -1

OpenStudy (anonymous):

(x-4)=16/-4 x=-4+4

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

How is it -1 ?

OpenStudy (anonymous):

-4(x-3)=16 -4x+12=16 -4x=4 x=-1

OpenStudy (anonymous):

Oh, Ok thanks

myininaya (myininaya):

Guys, please guide the user so they can answer the question themselves. Don't just give the answer and expect them to learn from just an answer. Thanks.

OpenStudy (anonymous):

@simplybeyoutiful please don't give out answers.

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

i was explaining it @Romero & you're welcome.

OpenStudy (anonymous):

You gave the answer. the asker even went beyond to ask again how you got that....and showing your work is not exactly explaining

myininaya (myininaya):

Honestly, there are some people who just come here looking for answers and we don't want to award those people by giving just the answer. After they get the answer, they don't care about your explanation. Luckly here, @CallMeKelly is a very good student and wants to learn. This will not always be the case.

OpenStudy (anonymous):

i showed them the steps. chill out.

OpenStudy (anonymous):

@simplybeyoutiful showed me the steps, now I understand why it was -1. Thats all I wanted to know

myininaya (myininaya):

Right I just want her to understand the rules. Thanks everyone.

OpenStudy (anonymous):

Thank you @CallMeKelly

OpenStudy (anonymous):

*high five*

OpenStudy (anonymous):

Your welcome